# envname-P envname#1

### v2

@@ 8 Future work directions
Appendix A Case Study: Logical Contradictions in Direct LLM Test Case Generation
@@
Appendix B Case Study: Contract Enforcement under Different Prompting Conditions
@@
Appendix C Case Study: Contract Adherence vs. Functional Correctness
@@
Appendix D Formalizing Contracts into SMT-LIB: Template and a Case Study
This section details the structure of the SMT-LIB template used by our PACT framework. SMT-LIB is a standardized, text-based language used to interface with SMT solvers. It provides a formal syntax for declaring variables, defining functions, and asserting logical formulas, allowing complex problems to be translated into a format that a solver can systematically analyze for satisfiability. Our framework leverages this language to translate nuanced, natural language contracts into a formal representation that can be reasoned about with logical precision.
@@ Appendix D Formalizing Contracts into SMT-LIB: Template and a Case Study
Figure 11 shows the ground-truth Python implementation for the HumanEval/11 task, which requires a function that takes two binary strings of equal length. The SMT-LIB formalization of these requirements is shown in Figure 12. The three assert statements in the Python code directly correspond to the three formal contracts defined in SMT-LIB:
  • C0 verifies that both inputs are strings, corresponding to the assertion assert isinstance(a, str) and isinstance(b, str).
  • C1 ensures their lengths are equal, corresponding to assert len(a) == len(b).
  • C2 checks that they are valid binary strings using a custom isBinaryString helper function, corresponding to assert set(a).issubset("0", "1") and set(b).issubset("0", "1").
Figure 10: The SMT-LIB template used for formalizing contracts.

### v3

@@ Appendix B Hidden Contracts and Contract-Violating Tests
Appendix C Case Study: Logical Contradictions in Direct LLM Test Case Generation
@@
Appendix D Case Study: Contract Enforcement under Different Prompting Conditions
@@
Appendix E Case Study: Contract Adherence vs. Functional Correctness
@@
Appendix F Formalizing Contracts into SMT-LIB: Template and a Case Study
This section details the structure of the SMT-LIB template used by the construction of ContractEval. SMT-LIB is a standardized, text-based language used to interface with SMT solvers. It provides a formal syntax for declaring variables, defining functions, and asserting logical formulas, allowing complex problems to be translated into a format that a solver can systematically analyze for satisfiability. Our framework leverages this language to translate nuanced, natural language contracts into a formal representation that can be reasoned about with logical precision.
Figure 12: The SMT-LIB template used for formalizing contracts.
@@ Appendix F Formalizing Contracts into SMT-LIB: Template and a Case Study
Figure 13 shows the ground-truth Python implementation for the HumanEval/11 task, which requires a function that takes two binary strings of equal length. The SMT-LIB formalization of these requirements is shown in Figure 14. The three assert statements in the Python code directly correspond to the three formal contracts defined in SMT-LIB: C0 verifies that both inputs are strings, corresponding to the assertion assert isinstance(a, str) and isinstance(b, str). C1 ensures their lengths are equal, corresponding to assert len(a) == len(b). C2 checks that they are valid binary strings using a custom isBinaryString helper function, corresponding to assert set(a).issubset("0", "1") and set(b).issubset("0", "1").
@@
Appendix G Full Result of Prompting Strategies
Table 4 reports the full metrics for the prompting variants discussed in Section 6.3. Beyond pass@1 and CSR, we include CodeBLEU and LLM-as-judge to capture syntactic and semantic alignment of the generated contract checks with the reference. Overall, CoT prompting is high-variance and often harmful: it can substantially degrade both functional correctness and contract satisfaction, and it catastrophically collapses for Phi-4 and Phi-4-plus (near-zero across all metrics). In contrast, the multi-turn strategy is markedly more stable. It typically recovers pass@1 relative to the single-turn EAS baseline (e.g., up to +4.2%p for Phi-4-plus), while keeping CSR competitive (within roughly 0–7%p of EAS across models). The auxiliary metrics largely track CSR: when CSR drops under a prompting variant, CodeBLEU and LLM-as-judge tend to drop as well, suggesting that performance changes are driven by missing or altered contract checks rather than superficial formatting differences.
Table 4: Comparison of additional prompting strategies. Multi-turn recovers pass@1 while maintaining competitive contract satisfaction to the standard EAS prompt.
Model Instruction Functional Contract Satisfaction
pass@1 ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
DeepSeek-R1 EAS 70.87% 52.66% 54.51% 81.49%
+ CoT 61.11% 35.00% 40.35% 62.46%
+Multi-Turn 72.96% 45.44% 45.75% 65.34%
gemma-3 EAS 69.56% 48.54% 48.88% 72.04%
+ CoT 70.46% 50.37% 50.68% 71.23%
+Multi-Turn 71.34% 49.26% 49.62% 72.45%
Phi-4 EAS 72.99% 50.73% 50.86% 75.63%
+ CoT 3.66% 1.91% 8.14% 13.28%
+Multi-Turn 71.88% 46.58% 48.44% 73.42%
Phi-4-plus EAS 73.21% 50.94% 51.04% 75.98%
+ CoT 0.00% 0.00% 2.92% 4.96%
+Multi-Turn 77.42% 48.52% 49.87% 75.01%
Qwen-3 EAS 69.65% 51.83% 53.30% 77.27%
+ CoT 71.13% 45.24% 50.01% 75.62%
+Multi-Turn 71.70% 49.23% 50.97% 76.98%
Appendix H Contract Evaluation on Functionally Correct Programs
Through ContractEval, our experiments show that prompting LLMs with contract information (e.g., CS and EAS) substantially improves contract satisfaction on ill-formed tests. However, contract satisfaction is only meaningful when the generated program is also functionally correct on well-formed inputs. If a program already fails on valid inputs, then its behavior on ill-formed inputs is ambiguous: it may “reject” an input simply because it crashes (e.g., due to unrelated bugs), rather than because it implements the intended contract checks. To disentangle true contract behavior from incidental failures, we analyze contract satisfaction conditioned on functional correctness in both Figure 15 and Table 5.
Figure 15: Contract satisfaction depending on the performance of functaional correctness.
Let $f_{i}\in[0,1]$ denote the functional test pass rate of a generated program for task $i$ (measured on well-formed tests), and let $c_{i}\in[0,1]$ denote its contract satisfaction (measured by CSR on ContractEval’s ill-formed tests). For a functionality threshold $t$, we consider the subset $\mathcal{S}_{t}=\{i\mid f_{i}\geq t\}$ and report: (i) the conditional mean contract satisfaction $\mathbb{E}[c\mid f\geq t]$, and (ii) the probability of perfect contract satisfaction $\Pr(c=1\mid f\geq t)$. Intuitively, increasing $t$ filters to progressively more functionally reliable programs, and reveals whether contract satisfaction improves as functional correctness improves.
Across models, we observe that contract satisfaction does not reliably track functional correctness. After removing the most functionally incorrect outputs (e.g., $t\geq 0.25$), $\mathbb{E}[c\mid f\geq t]$ becomes largely flat: making a program more functionally correct does not automatically make it more contract-satisfying. Even among fully functionally correct programs ($f=1$), a substantial fraction still fails to reject contract-violating inputs. For example, under our contract-aware prompting setting, only about half of functionally correct programs achieve perfect contract satisfaction ($\Pr(c=1\mid f=1)$ is roughly in the 0.51–0.59 range across models), indicating that many solutions that “solve the task” still omit essential validity checks.
This conditional analysis supports the central premise of ContractEval: functional correctness on well-formed inputs is insufficient to characterize robustness. Contract satisfaction remains a distinct and non-trivial dimension of code quality, and ContractEval’s ill-formed tests expose failures that are invisible under pass@k-style evaluation.
Table 5: Contract satisfaction evaluation on functionally correct programs.
Model Instruction Correct Functions Contract Satisfaction
Ratio ($\uparrow$) CSR ($\uparrow$) CodeBLEU ($\uparrow$) LLM-as-judge ($\uparrow$)
DeepSeek-R1 CS 63.19% 32.89% 35.96% 55.38%
EAS 59.34% 56.75% 55.01% 83.01%
gemma-3 CS 64.84% 26.53% 31.98% 44.61%
EAS 57.14% 54.29% 54.20% 79.99%
Phi-4 CS 61.26% 52.01% 52.25% 78.34%
EAS 61.81% 61.78% 56.33% 84.21%
Phi-4-plus CS 59.89% 49.58% 48.11% 71.89%
EAS 62.91% 61.82% 55.31% 83.31%
Qwen-3 CS 60.99% 30.05% 39.93% 55.83%
EAS 59.34% 58.76% 57.79% 85.18%